### v2

@@ 3 Coderivatives of the second-order cone
$\begin{array}[]{ll}\langle{\boldsymbol{r}},(1;-{\boldsymbol{w}})\rangle=-\langle{\boldsymbol{p}}+{\boldsymbol{q}},(1;-{\boldsymbol{w}})\rangle=-\dot{p}+\langle\bar{\boldsymbol{p}},{\boldsymbol{w}}\rangle-\dot{q}+\langle\bar{\boldsymbol{q}},{\boldsymbol{w}}\rangle\\ =-\dot{p}+\langle\bar{\boldsymbol{p}},\frac{\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{p}}\|}\rangle-\dot{q}+\langle\bar{\boldsymbol{q}},\frac{\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{p}}\|}\rangle=-\dot{p}+\|\bar{\boldsymbol{p}}\|+\frac{1}{\|\bar{\boldsymbol{p}}\|}(\dot{q}\dot{p}+\langle\bar{\boldsymbol{q}},\bar{\boldsymbol{p}}\rangle)=2\|\bar{\boldsymbol{p}}\|+\frac{\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle}{\|\bar{\boldsymbol{p}}\|}>0.\end{array}$ (3.12)
@@
which, together with (3.12) and the fact that $\varrho\in(0,1]$, implies ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ from (3.11).
@@
${\boldsymbol{s}}-{\boldsymbol{r}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{s}}),\quad\forall\,{\boldsymbol{r}}\in{\mathds{R}}^{1+r},$ (3.13)
@@
$\begin{array}[]{ll}\langle{\boldsymbol{r}},(1;{\boldsymbol{w}})\rangle=-\langle{\boldsymbol{p}}+{\boldsymbol{q}},(1;{\boldsymbol{w}})\rangle=-\dot{p}-\langle\bar{\boldsymbol{p}},{\boldsymbol{w}}\rangle-\dot{q}-\langle\bar{\boldsymbol{q}},{\boldsymbol{w}}\rangle\\ =-\dot{p}-\langle\bar{\boldsymbol{p}},-\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle-\dot{q}-\langle\bar{\boldsymbol{q}},-\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle=-\dot{p}+\langle\bar{\boldsymbol{p}},\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle-\dot{q}+\|\bar{\boldsymbol{q}}\|=\frac{\langle{\boldsymbol{q}},{\boldsymbol{p}}\rangle}{\|\bar{\boldsymbol{q}}\|}+2\|\bar{\boldsymbol{q}}\|>0.\end{array}$ (3.14)
@@
This, together with (3.13), (3.14), and the fact that $\varrho\in(0,1]$, implies that ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ holds.
(6e) Since ${\boldsymbol{u}}={\boldsymbol{v}}={\boldsymbol{0}}$, by Part (6) of Lemma 3.1 and (3.1) one has for any ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$, it holds ${\boldsymbol{r}}\in{\mathcal{D}}^{\ast}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}})({\boldsymbol{r}})$. This is equivalent, by Lemma 2.3, to ${\boldsymbol{0}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{r}})$ for all ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$. This completes the proof of the lemma. ∎
@@ 4 Equivalence of the Aubin property and the strong regularity
This section establishes the main result of this paper. The main step is to show that the strong second-order sufficient condition in Definition 4.1 holds. We will reformulate the Mordukhovich criterion for the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ to the necessary condition in Corollary 4.1. Moreover, by specifying a basis for the affine hull of the critical cone in Corollary 4.2, we can put our discussion in a space with a lower dimension (Lemma 4.3). Finally, by extensively utilizing Corollary 4.1 and Lemma 2.1, we obtain the strong second-order sufficient condition by Propositions 4.2 and 4.3, in which successive dimension reduction is involved.
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to (1.1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier. Note that any index $j\in\{1,\ldots,J\}$ belongs to one and only one of the following $6$ sets:
$\left\{\begin{array}[]{ll}J_{1}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}\in{\sf int}{\mathcal{Q}}_{j}\right\},\\[2.84526pt] J_{2}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})\in{\sf int}{\mathcal{Q}}_{j},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\},\\[2.84526pt] J_{3}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*}),({\bm{\mu}}^{*})^{j}\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\}\right\},\\[2.84526pt] J_{4}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\}\right\},\\[2.84526pt] J_{5}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\},\\[2.84526pt] J_{6}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\}.\end{array}\right.$ (4.1)
@@
${\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}))=\begin{cases}{\mathds{R}}^{1+r_{j}},&\mbox{ if }j\in J_{2},\\ {\mathcal{Q}}_{j},&\mbox{ if }j\in J_{1,4,6},\\ \{{\boldsymbol{v}}\in{\mathds{R}}^{1+r_{j}}\mid\langle\bar{\boldsymbol{v}},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}\dot{g}^{j}({\boldsymbol{x}}^{*})\leq 0\},&\mbox{ if }j\in J_{3,5}.\end{cases}$ (4.2)
@@
${\mathcal{C}}({\boldsymbol{x}}^{*})=\left\{{\boldsymbol{d}}\in{\mathds{R}}^{n}\middle|\begin{array}[]{ll}{\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},&\forall\,j\in J_{1},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*})),&\forall\,j\in J_{5,6},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathds{R}}_{+}(({\dot{\mu}^{*}})^{j};-({\bar{\bm{\mu}}^{*}})^{j}),&\forall\,j\in J_{4},\\ \langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0,&\forall\,j\in J_{3}\end{array}\right\}.$ (4.3)
@@
${\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))=\left\{{\boldsymbol{d}}\in{\mathds{R}}^{n}\middle|\begin{array}[]{ll}{\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},&\forall\,j\in J_{1},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathds{R}}(({\dot{\mu}^{*}})^{j};-({\bar{\bm{\mu}}^{*}})^{j}),&\forall\,j\in J_{4},\\ \langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0,&\forall\,j\in J_{3}\end{array}\right\}.$ (4.4)
In the following definition, the second-order and strong second-order sufficient conditions for (1.1) come from [6, Theorems 29 & 30].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point of the nonlinear SOCP (1.1). We say that the second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists an associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
$\langle{\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\rangle>0,\quad\forall\,{\boldsymbol{d}}\in{\mathcal{C}}({\boldsymbol{x}}^{*})\setminus\{{\boldsymbol{0}}\},$ (4.5)
@@
${\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\right>>0,\quad\forall\,{\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}.$ (4.6)
@@
${\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\right>\geq 0,\quad\forall\,{\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*})).$ (4.7)
@@
We will show that $\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\neq{\boldsymbol{0}}$ for all ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}$ (Corollary 4.1) if the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ holds. Consequently, we devote our main effort to proving that the strong second-order necessary condition (4.7) holds to obtain Proposition 4.3. In [3, Theorem 4.2], by assuming that the variational sufficiency for local optimality [31] holds, the authors established the equivalence between the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ and the strong regularity of (1.6) for stationary points. It can be observed from [31, Theorem 3 & Example 1] that for the conventional nonlinear programming, i.e., a special case of (1.2) with $r_{1}=\cdots=r_{J}=0$, the variational sufficiency for local optimality is a sufficient condition for the strong second-order necessary condition (4.7). Therefore, making such an assumption is very close to directly assuming the strong second-order sufficient condition (4.6). In contrast, we will derive the strong second-order sufficient condition solely from the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ without assuming the variational sufficiency as in [31].

### v3

@@ 3 Coderivatives of the second-order cone
(24) $\begin{array}[]{ll}\langle{\boldsymbol{r}},(1;-{\boldsymbol{w}})\rangle=-\langle{\boldsymbol{p}}+{\boldsymbol{q}},(1;-{\boldsymbol{w}})\rangle=-\dot{p}+\langle\bar{\boldsymbol{p}},{\boldsymbol{w}}\rangle-\dot{q}+\langle\bar{\boldsymbol{q}},{\boldsymbol{w}}\rangle\\ =-\dot{p}+\langle\bar{\boldsymbol{p}},\frac{\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{p}}\|}\rangle-\dot{q}+\langle\bar{\boldsymbol{q}},\frac{\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{p}}\|}\rangle=-\dot{p}+\|\bar{\boldsymbol{p}}\|+\frac{1}{\|\bar{\boldsymbol{p}}\|}(\dot{q}\dot{p}+\langle\bar{\boldsymbol{q}},\bar{\boldsymbol{p}}\rangle)=2\|\bar{\boldsymbol{p}}\|+\frac{\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle}{\|\bar{\boldsymbol{p}}\|}>0.\end{array}$
@@
which, together with (24) and the fact that $\varrho\in(0,1]$, implies ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ from (23).
@@
(25) ${\boldsymbol{s}}-{\boldsymbol{r}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{s}}),\quad\forall\,{\boldsymbol{r}}\in{\mathds{R}}^{1+r},$
@@
(26) $\begin{array}[]{ll}\langle{\boldsymbol{r}},(1;{\boldsymbol{w}})\rangle=-\langle{\boldsymbol{p}}+{\boldsymbol{q}},(1;{\boldsymbol{w}})\rangle=-\dot{p}-\langle\bar{\boldsymbol{p}},{\boldsymbol{w}}\rangle-\dot{q}-\langle\bar{\boldsymbol{q}},{\boldsymbol{w}}\rangle\\ =-\dot{p}-\langle\bar{\boldsymbol{p}},-\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle-\dot{q}-\langle\bar{\boldsymbol{q}},-\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle=-\dot{p}+\langle\bar{\boldsymbol{p}},\frac{\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{q}}\|}\rangle-\dot{q}+\|\bar{\boldsymbol{q}}\|=\frac{\langle{\boldsymbol{q}},{\boldsymbol{p}}\rangle}{\|\bar{\boldsymbol{q}}\|}+2\|\bar{\boldsymbol{q}}\|>0.\end{array}$
@@
This, together with (25), (26), and the fact that $\varrho\in(0,1]$, implies that ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ holds.
(6e) Since ${\boldsymbol{u}}={\boldsymbol{v}}={\boldsymbol{0}}$, by Part (6) of Lemma 3.1 and (13) one has for any ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$, it holds ${\boldsymbol{r}}\in{\mathcal{D}}^{\ast}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}})({\boldsymbol{r}})$. This is equivalent, by Lemma 2.4, to ${\boldsymbol{0}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{r}})$ for all ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$. This completes the proof of the lemma.
@@ 4 Equivalence of the Aubin property and the strong regularity
This section establishes the main result of this paper. The main step is to show that the strong second-order sufficient condition in Definition 4.1 holds. We will reformulate the Mordukhovich criterion for the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ to the necessary condition in Corollary 4.9. Moreover, by specifying a basis for the affine hull of the critical cone in Corollary 4.11, we can put our discussion in a space with a lower dimension (Lemma 4.13). Finally, by extensively utilizing Corollary 4.9 and Lemma 2.1, we obtain the strong second-order sufficient condition by Propositions 4.15 and 4.17, in which successive dimension reduction is involved.
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to (1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier. Note that any index $j\in\{1,\ldots,J\}$ belongs to one and only one of the following $6$ sets:
(27) $\left\{\begin{array}[]{ll}J_{1}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}\in{\sf int}{\mathcal{Q}}_{j}\right\},\\[2.84526pt] J_{2}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})\in{\sf int}{\mathcal{Q}}_{j},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\},\\[2.84526pt] J_{3}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*}),({\bm{\mu}}^{*})^{j}\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\}\right\},\\[2.84526pt] J_{4}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\}\right\},\\[2.84526pt] J_{5}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{{\mathcal{Q}}_{j}}\setminus\{{\boldsymbol{0}}\},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\},\\[2.84526pt] J_{6}:=\left\{j\mid g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}},({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}\right\}.\end{array}\right.$
@@
(28) ${\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}))=\begin{cases}{\mathds{R}}^{1+r_{j}},&\mbox{ if }j\in J_{2},\\ {\mathcal{Q}}_{j},&\mbox{ if }j\in J_{1,4,6},\\ \{{\boldsymbol{v}}\in{\mathds{R}}^{1+r_{j}}\mid\langle\bar{\boldsymbol{v}},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}\dot{g}^{j}({\boldsymbol{x}}^{*})\leq 0\},&\mbox{ if }j\in J_{3,5}.\end{cases}$
@@
(29) ${\mathcal{C}}({\boldsymbol{x}}^{*})=\left\{{\boldsymbol{d}}\in{\mathds{R}}^{n}\middle|\begin{array}[]{ll}{\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},&\forall\,j\in J_{1},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*})),&\forall\,j\in J_{5,6},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathds{R}}_{+}(({\dot{\mu}^{*}})^{j};-({\bar{\bm{\mu}}^{*}})^{j}),&\forall\,j\in J_{4},\\ \langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0,&\forall\,j\in J_{3}\end{array}\right\}.$
@@
(30) ${\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))=\left\{{\boldsymbol{d}}\in{\mathds{R}}^{n}\middle|\begin{array}[]{ll}{\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}},&\forall\,j\in J_{1},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\in{\mathds{R}}(({\dot{\mu}^{*}})^{j};-({\bar{\bm{\mu}}^{*}})^{j}),&\forall\,j\in J_{4},\\ \langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0,&\forall\,j\in J_{3}\end{array}\right\}.$
In the following definition, the second-order and strong second-order sufficient conditions for (1) come from [6, Theorems 29 & 30].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point of the nonlinear SOCP (1). We say that the second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists an associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
(31) $\langle{\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\rangle>0,\quad\forall\,{\boldsymbol{d}}\in{\mathcal{C}}({\boldsymbol{x}}^{*})\setminus\{{\boldsymbol{0}}\},$
@@
(32) ${\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\right>>0,\quad\forall\,{\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}.$
@@
(33) ${\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\right>\geq 0,\quad\forall\,{\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*})).$
@@
We will show that $\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\neq{\boldsymbol{0}}$ for all ${\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}$ (Corollary 4.9) if the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ holds. Consequently, we devote our main effort to proving that the strong second-order necessary condition (33) holds to obtain Proposition 4.17. In [3, Theorem 4.2], by assuming that the variational sufficiency for local optimality [31] holds, the authors established the equivalence between the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ and the strong regularity of (6) for stationary points. It can be observed from [31, Theorem 3 & Example 1] that for the conventional nonlinear programming, i.e., a special case of (2) with $r_{1}=\cdots=r_{J}=0$, the variational sufficiency for local optimality is a sufficient condition for the strong second-order necessary condition (33). Therefore, making such an assumption is very close to directly assuming the strong second-order sufficient condition (32). In contrast, we will derive the strong second-order sufficient condition solely from the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ without assuming the variational sufficiency as in [31].